OpenStudy (anonymous):

Three times the sum of a number and seven is two less than five times the number.

OpenStudy (anonymous):

Write an algebraic equation for the following.

OpenStudy (anonymous):

do u have any options to choose from?

OpenStudy (anonymous):

yes, A. –4 B. 9.6 C. 8 D. 16

OpenStudy (anonymous):

whats a?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

you wouldnt be able to solve this equation bcuz there is no value for the varible

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

were those numbers you gave me, were they answers to choose from or values of the varible?

OpenStudy (anonymous):

no these numbers A. 3x + 7 = 5x – 2 B. 3(x + 7) = 5x – 2 C. 3(x + 7) = 2 – 5x D. 3x + 7 = 2 – 5x

OpenStudy (anonymous):

are all those questions marks supposed to be a - sign

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then what are they

OpenStudy (anonymous):

A. 3x + 7 = 5x – 2 B. 3(x + 7) = 5x – 2 C. 3(x + 7) = 2 – 5x D. 3x + 7 = 2 – 5x

OpenStudy (anonymous):

\[ A. 3x + 7 = 5x – 2 B. 3(x + 7) = 5x – 2 C. 3(x + 7) = 2 – 5x D. 3x + 7 = 2 – 5x\]

OpenStudy (anonymous):

is the answer for c 3(x+7) = 2-5x

OpenStudy (anonymous):

is it yes or no